### v1

# A -ray-emitting blazar B3 1239+376 at = 3.82 identified in a multi-wavelength context

Abstract
Among thousands of extragalactic -ray emitters, only a handful of distant ( 3) sources are detected, yet they are cruial probes shedding light on the cosmic evolution of jets of active galactic nuclei and the initial phase of mass growth of supermassive black holes. Here, we report on a multi-band study of a radio quasar B3 1239+376 with = 3.82. By analyzing the Fermi-LAT data, a significant (globally 7.7) -ray source in its direction, with an estimated association probability of 0.91, is observed in a half-year period of 2025. The analyses also reveal the emergence of co-spatial -ray residues in prior epochs. Moreover, the -ray and infrared light curves are likely correlated, particularly, the two emissions climb to the peaking values at the same time. The temporal coincidence establishes a firm association relationship between the -ray source and the quasar. Therefore, B3 1239+376 is proposed as the third most distant -ray-emitting blazar to date. Benefiting from the multi-wavelength observations, broadband spectral energy distributions in different flux states are drawn and reproduced by the classic one-zone leptonic radiation model to investigate the jet properties. Considering the recent brightening in rays, prompt follow-up observations are encouraged, especially radio interferometry observations which may catch the potential ejection of a new jet blob.
I Introduction
Active galactic nuclei (AGNs), are one of the most luminous sources of electromagnetic emissions in the universe, powered by the central supermassive black holes (SMBHs, Antonucci 1993; Urry & Padovani 1995). In the past 10 years, more than 400 quasars with z 5.7 have been discovered (Fan etĀ al., 2023). The existence of these objects suggests a prompt and eļ¬cient mass growth for the SMBHs, challenging the scenarios of SMBH formation in the early universe (Inayoshi etĀ al., 2020). The situation becomes more interesting when the high- AGNs with strong jets (i.e., jetted AGN) are detected (e.g., An & Romani, 2018; Belladitta etĀ al., 2020; Sbarrato etĀ al., 2022; BaƱados etĀ al., 2025). In this case, since the presence of relativistic jets is believed to relate with highly spinning accreting black holes (e.g., Tchekhovskoy etĀ al., 2011) that a high value for the radiation eļ¬ciency is expected (, Thorne 1974), a longer growth time of these sources is needed than that of the non-jetted AGNs.
Blazars, including flat-spectrum radio quasars (FSRQs) and BL Lacertae objects (BL Lacs), are a subclass of AGNs that their strong jets are aligned with the line of sight and hence the beamed non-thermal jet emissions are overwhelming (Blandford etĀ al., 2019). A distinguished feature of the blazar radiation is the violent variability across diverse timescales (e.g., Ulrich etĀ al., 1997). The broadband jet emissions span from radio to very-high-energy -ray domains, characterized by a universal two-bump spectral shape in logālog plot. One bump is widely accepted from synchrotron emission, while the other one peaking in rays, is usually explained as inverse Compton (IC) by a population of relativistic electrons, which are also responsible for the synchrotron emission, scattering of soft photons from either inside (the synchrotron self-Compton, or SSC, Maraschi etĀ al. 1992) and/or outside (external Compton, or EC, Dermer & Schlickeiser 1993; Sikora etĀ al. 1994; BÅażejowski etĀ al. 2000) the jet. Alternatively, hadronic radiation models are adopted to describe the spatial and temporal coincidences between the incoming neutrino and the flaring blazars (e.g., IceCube Collaboration etĀ al., 2018; KM3NeT Collaboration etĀ al., 2025).
In the -ray sky, the majority of the extragalactic sources are blazars (Ballet et al., 2023). Although the most distant blazar detected so far is with redshift as high as = 7 (Bañados et al., 2025), the redshift record of -ray-emitting blazars is 4.72 right now (i.e., B3 1428+422, Liao et al. 2018; Kreter et al. 2020; Gokus et al. 2025). Actually, there are only two blazars with -ray detections beyond redshift of 4, the other one is GB 1508+5714 ( = 4.3) (Ackermann et al., 2017; Liao et al., 2020; Gokus et al., 2024). The -ray radiations bring key information of the jet properties (Madejski & Sikora, 2016). In particular, -ray photons from high- blazars, are a powerful tool for probing the extragalactic background light (EBL, Fermi-LAT Collaboration et al. 2018). Since these sources are spectrally soft (Ackermann et al., 2017; Li et al., 2018) and the angular resolution of Fermi-LAT for the sub-GeV photons is worse than that of GeV photons, the major obstacle to detecting -ray signals from high- blazars is the limited localization. Catching correlated multi-wavelength activities between the -ray emission and those in other windows of electromagnetic radiation acts as a decisive evidence to determine the association relationship (e.g., Liao et al., 2019).
B3 1239+376 is first known as a radio source observed by the Bologna Northern Cross Telescope (Colla etĀ al., 1973). It possesses a flat radio spectrum between 1.4 and 5 GHz (Taylor etĀ al., 1996) and is recognized as a high- quasar ( = 3.82, Vermeulen etĀ al. 1996), which makes it one of the brightest radio quasars in the early universe (i.e., 0.5 Jy). Although B3 1239+376 is not listed in the Roma-BZCat (hereafter BZCAT, Massaro etĀ al. 2009), the super-luminous motion of 5 has been detected (Britzen etĀ al., 2008). Meanwhile, its radio-loudness (RL=, Kellermann etĀ al. 1989) is calculated as high as 25,000, suggesting a blazar nature. In the X-ray perspective, a hard X-ray spectrum (i.e., = 1.5) is suggested, further supporting it as a blazar (McKeough etĀ al., 2016). In this study, multiwavelength data of B3 1239+376 are collected and analyzed, and its broadband temporal properties are reported (Section II), along with some discussions (Section III). Here we take a CDM cosmology with , , and (Planck Collaboration etĀ al., 2014).
II Data Analysis and Results
II.1 Fermi-LAT Data
We collected Fermi-LAT Pass 8 SOURCE data, with an energy range between 100 MeV and 500 GeV. The data were taken during the period from 2008 August 4 to 2025 August 4 (i.e., MJD 54683 - 60892), at a time length of 17-year. During the analysis, Fermitools software version 2.2.0 as well as Fermitools-data version 0.18 were adopted. A zenith angle cut (i.e., 90ā) along with the recommended quality-filter cuts (i.e., DATA_QUAL==1 && LAT_CONFIG==1) were included in the data filtering procedures. In the Unbinned likelihood analyses, in which -ray flux and spectrum were extracted by the gtlike task, the test statistic (TS = -2ln (/), Mattox etĀ al. 1996) was used to estimate the the significance of the -ray sources. In the definition, and denote the maximum likelihood values for the models with and without the target -ray source, respectively. Since the target is not embraced in an incremental version (data release 4) of the fourth Fermi-LAT catalog (4FGL-DR4, Ballet etĀ al. 2023), a test -ray sources at the radio position of B3 1239+376 was added in the source model. During the likelihood analysis, a region of interest (ROI) of 10ā centered on the target was considered. Parameters of the 4FGL-DR4 sources within the ROI, and those for the diffuse emission templates were set free, while others were frozen as the default values. In the temporal analysis, if detections of the background sources are marginal (i.e., TS 10), they were removed from the source model. In this case, for the target, 95% confidence level (CL) upper limits were calculated by the pyLikelihood UpperLimits tool, to replace the flux estimations.
Since 4FGL-DR4 is based on the first 14-yr Fermi-LAT survey, an analysis for the additional 3-year data were performed. Interestingly, an excess in this direction emerged, with a TS value of 47 (locally 5.9). Therefore, a two-month time bin -ray light curve, was extracted to seek any potential activities recently. As shown in Figure 1, the background emission was dominant until the recent half-year period (i.e., MJD 60710 - 60892). Subsequent analysis aiming this specific half-year data revealed a significant residue (TS = 80, locally 8.2) towards B3 1239+376, see Figure 2. Here there are 34 trials (i.e., a selected half-year period for the entire 17-year dataset) and hence a global significance, after correction of the trial factor, was estimated as 7.7, suggesting that the detection is statistically significant. Assuming a simple power-law spectral template (i.e. , where is the photon index), the analysis yields a photon flux of ph . The source is spectrally soft, with an estimated photon index of . By comparison, the fit of the prior 2.5-year data only provided an upper limit (TS = 8), ph . Therefore, a strong variation with a 5-fold flux increase is indicated.
Considering the limited angular resolution of Fermi-LAT, the possible influences from the background sources were checked. The closest background source is 4FGL J1243.2+3627, about 0.9ā away, however, it possesses a rather different spectral behavior (, Ballet etĀ al. 2023) than our target. We noted that the brightest background source in the ROI (i.e., 4FGL J1310.5+3221), which is 7.7ā from the ROI center, underwent a strong outburst in year 2025111https://fermi.gsfc.nasa.gov/ssc/data/access/lat/LightCurveRepository/source.html?source_name=4FGL_J1310.5+3221. Because the 95% CL containment angle of Fermi-LAT for 500 MeV photon is 222https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm, by selecting events with energies beyond 500Ā MeV, a further analysis was carried out. In this case, the contamination from the bright source could be significantly diminished. The detection remains to be significant, with a TS value of 39 which corresponds to a global significance of 4.7. Therefore, the new -ray source is likely intrinsic than artificially caused by the background sources. The obtained photon flux and index from the 500Ā MeV data analysis are consistent with that from the full data energy range.
Localization analysis of the -ray source was followed. The optimized location was constrained as R.A. 190ā.6411 and decl. 37ā.2587, with a 95% CL uncertainty radius of 0.22ā. Corresponding angular separation between that and the radio position of B3 1239+376 is 0.11ā, and hence it falls into the -ray localization area, see Figure 2. When low energy threshold of the photons was set to 500Ā MeV, the best location is R.A. 190ā.4602 and decl. 37ā.2475. At this time, the angular separation (0.11ā) is still smaller than the a 95% CL uncertainty radius (0.21ā). Other potential low-energy association candidates were searched. Since blazars are the dominant population of the extragalactic sky, the blazar candidate catalogs were firstly looked into. No co-spatial sources listed in the BZCAT list (Massaro etĀ al., 2009) were found. Moreover, the result holds when the WIBRaLS2 catalog (DāAbrusco etĀ al., 2019) was investigated. Additionally, low-energy counterparts for -ray source have been searched in the high-frequency radio survey as well as from the VLBI observations (Abdollahi etĀ al., 2020). B3 1239+376 is the only one embraced by the CRATES catalog (Healey etĀ al., 2007) as well as the radio fundamental catalog (Petrov & Kovalev, 2025) in this direction. Adopting the Bayesian association method together with the corresponding prior probability value for CRATES catalog (0.33, Abdo etĀ al. 2010), the association probability was calculated as 0.91, above the threshold value of 0.8. Hence, B3 1239+376 is suggested to associate with the -ray source. Nevertheless, considering the relatively large localization uncertainty in the rays, detections of multiwavelength correlated emissions are needed to establish a firm association relationship.
To provide a global picture of the temporal behavior of the -ray source, a 1-year time bin light curve for the entire 17-year dataset was built, as shown in Figure 3. Adopting the āvariability indexā test (Nolan etĀ al., 2012), the variability is suggested to be significant (). The recent brightening (i.e., the last time bin) stands out, and it has been picked up by the Bayesian block approach (Scargle etĀ al., 2013) indicative of a flare event, which is consistent with the results mentioned above. At the initial 9-year Fermi-LAT observation, the target was in a low flux state. The corresponding analysis yielded a marginal detection (TS =12), with a photon flux constrained to ph by setting the photon index as 2.6. It is noteworthy that there are tentative detections (TS 15) in two yearly time bins around year 2017 and 2020, respectively. Then a 4-year long quiescent state was followed, from which an upper limit of ph was obtained.
Detailed investigations on the potential signals from year 2017 to 2020 were carried out. In the first step, a 1-month time bin light curve was generated, to identify the exact time range of the signals. Two epochs were selected, in which one is a 6-month period between MJD 58060 and 58242, and the other one lasting for 10 months is from MJD 58727 to 58969. The corresponding obtained TS values are 15 and 17, respectively. Since analyzing Fermi-LAT data in neither of the epochs alone yield significant detections (i.e., TS 25), a joint analysis of the two epochs was then performed. The joint analysis returned an increased TS value (TS = 31) and the enhanced statistics allowed us to localize the residue. The optimized location was gave as R.A. 190ā.642 and decl. 37ā.4649, along with a 95% CL uncertainty radius of 0.21ā. As shown in Figure 2, B3 1239+376 is included within the location error radius, since the angular separation is 0.15ā under this circumstance. The photon flux and the spectral information were also extracted. The former was obtained as ph , while the photon index was constrained as that is in agreement with the result from the flare in year 2025. Both the spatial and spectral information extracted in the joint analysis suggest the potential activities then likely also related to B3 1239+376.
II.2 Chandra X-ray Observatory data
B3 1239+376 was observed twice by the Advanced CCD Imaging Spectrometer (ACIS, Garmire etĀ al. 2003), on 2007 Mar. 10 (ACIS-7, ObsID: 7871, PI.: C. Cheung) and 2022 Dec. 31 (ACIS-67, ObsID 26906; PI: E. Meyer). The corresponding exposure times are 5.2 and 20.1 ks, respectively. The observations were performed on the pointing mode, events were recorded in the Faint telemetry format and with the Timed Exposure mode. Data reductions were carried out using the Chandra Interactive Analysis of Observations software package (Fruscione etĀ al., 2006) v4.17 with CALDB version 4.12. Initially, the standard level 1 event lists were reprocessed into the level 2 data products by the chandra_repro script, in which standard grade, status, and good time filters, as well as VFAINT background cleaning were applied. Meanwhile, the routine deflare was utilized to remove background flaring periods. Pileup was estimated using PIMMS version 4.15111https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl. Since B3 1239+376 had % pileup over the 0.5ā7.0 keV energy band, pileup was ignored for this study.
The specextract task was used in the X-ray spectral analysis. The target events were extracted from a circular region with a radius of centered at the position of B3 1239+376, while for background photons nearby circles with a larger radius in the blank region were considered. The numbers of observed net photons of the observation in year 2007 and 2022 are 117 and 198, respectively. We grouped the spectra to have at least five counts per bin. Then the XPEC tool (Arnaud 1996, version 12.14.0h) was used to archive the spectral fitting procedure, in which the the energy range was set as 0.5 ā 7Ā keV and the -statistic (Cash, 1979) was adopted. The spectral model was assumed as a power-law distribution absorbed by the Galactic column density along the line of sight ( = , HI4PI Collaboration etĀ al. 2016). The data together with the optimized models were shown in Figure 4. The analysis on the observation in year 2007 gave an unabsorbed flux of (-Statistic/d.o.f., 12.3/17), along with a hard spectrum . While the results on the other observation indicates a possibly fainter emission, (-Statistic/d.o.f., 26.3/33). The X-ray emission in year 2022 is confirmed to be spectrally hard, .
II.3 SDSS spectrum
During the fit of the SDSS spectrum, effects of galactic extinction was removed firstly, and the continuum was described by a power law and polynomial model plus the Fe ii template (Vestergaard & Wilkes, 2001; Vestergaard & Peterson, 2006). Uncertainties were derived from 1000 mock spectra, generated by adding Gaussian noise consistent with the measurement errors. The standard deviation of the best-fit parameters across these realizations provided the uncertainty for each spectral quantity. As shown in Figure 5, several emission lines are distinct, especially the Ly and the Cāiv lines. The redshift is constrained as =3.82 0.01, consistent with the value list in SDSS quasar catalog DR16 (Wu & Shen, 2022). For the Ly emission line region, we used two components to fit Ly line and one component to fit NV line. The is obtained as , while the is . Taking the bolometric correction factor = 3.81 (Shen etĀ al., 2011), the bolometric disk luminosity is suggested as high as . As for the Cāiv line fitting, three Gaussian components, including one narrow component and two broad components, were adopted. We estimated the as . However, note that the Cāiv line is known to be biased by strong outflows, and hence it should be cautious for virial black hole mass estimations (Dix etĀ al., 2023). Future near-infrared spectroscopy is needed to give a robust BH mass estimation of B3 1239+376 (Burke etĀ al., 2024).
II.4 Infrared data
B3 1239+376 is included in the AllWISE Source Catalog (Wright etĀ al., 2010; WISE Team, 2019), named as WISEA J124209.81+372006.0. The W1 (m) and W2(m) magnitudes listed in the catalog are 16.22 0.06 and 15.90 0.15, respectively. The detections in W3 (m) and W4(m) bands are not statistically significant. Long-term infrared light curves in W1 and W2 bands were extracted, based on the time-resolved WISE/NEOWISE co-added imagesĀ (Meisner etĀ al., 2018), due to the faint infrared radiations. During the analyses, point spread function photometry was carried out, targeting on the the difference images that the first epoch observations were treated as reference images (Jiang etĀ al., 2021). Along with the long-term -ray light curve, the infrared light curves are also shown in Figure 3. A -test with a constant flux level as the null hypothesis was utilized to investigate the significance of the infrared variations. The minimum reduced values obtained by varying the constant flux levels in W1 and W2 light curves are 7.2 and 3.9 (d.o.f. = 23), respectively. Therefore, the infrared variability is suggested to be significant (). The variation tendency in the two bands are similar. The variability amplitude in W2 band appears to be higher than that in the W1 band, but the former also has larger measure uncertainties. At the start of the WISE survey, B3 1239+376 was in the quiescent state, after year 2014, a flux rise was apparent lasting about 1500 days. The detected flux maxima were at MJD 58260, then a flux decline was followed. The emissions bounced back with a peak at MJD 58832. There is a minor flux activity later around MJD 59562.
Although the WISE mission was ended at 31th July 2024 (i.e., MJD 60522), a new infrared space telescope, named as Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer (SPHEREx, Bock et al. 2025), has been successfully launched in Mar. 2025. It performs a 0.75-5 µm spectroscopy survey over the entire sky. Based on its first reprocessed quick release products (SPHEREx Team, 2025), one spectrum on B3 1239+376 observed from MJD 60800 to MJD 60829 is obtained. To compare with the WISE measurements, we convolved the SPHEREx spectrum with the WISE W1 and W2 filters transmission curves222https://svo2.cab.inta-csic.es/svo/theory/fps3/index.php. The obtained SPHEREx W1 and W2 flux densities are given as (0.259 0.008) mJy and (0.278 0.32) mJy, respectively, see Figure 3. Interestingly, these values are unprecedentedly high, roughly 2-3 times of majority of the WISE estimations.
II.5 Radio data
B3 1239+376 is a well studied radio source. It was monitored by RATAN-600 (Parijskij, 1993), in 12 epochs between MJD 57836 and 59026, at multi-bands from 1.2Ā GHz to 22Ā GHz (Sotnikova etĀ al., 2021). In addition, seven snapshots at K-band (24Ā GHz) of Very Long Baseline Array (VLBA) observations with submilliarcsecond resolution were carried out (de Witt etĀ al., 2023). Images from the Very Large Array Sky Survey (VLASS, Lacy etĀ al. 2020) were also looked into, from which radio fluxes at 3Ā GHz were extracted by modelling the source with a Gaussian fit on the image plane using the Common Astronomy Software Applications. The VLASS flux density in the first epoch was obtained as 612 9 mJy, which is consistent with the value (i. e., 608 3 mJy) in epoch 1 quick look catalog of VLASS (Gordon etĀ al., 2020). The radio data are also plotted in Figure 3. The RATAN-600 multi-bands light curves exhibit a general behavior of flux decay in long-timescale, which is supported by the results of the VLASS observations. Two components, including a compact core as well as a parsec-scale jet, are resolved in the VLBA images, interestingly, from which different temporal behaviors are revealed. For the core, the K-band flux density is 0.099 Jy at MJD 57807 and then reaches to 0.153 Jy at MJD 58180, with an average relative uncertainty of 9%, indicating a likely flux increase. On the other hand, no significant variation of the jet flux density is found.
II.6 Implications from the multi-wavelength observations
Through a detailed investigation of the multi-wavelength properties of B3 1239+376, its nature as a high- -ray-emitting blazar is revealed. In addition to its typical blazar behavior recorded in literature, for instance, the extremely high radio loudness, super-luminous motion and hard X-ray spectrum (Britzen etĀ al., 2008; McKeough etĀ al., 2016), in this study, -ray signals detected by Fermi-LAT toward B3 1239+376 are reported. In year 2025, a robust (7.7 in global) -ray source there emerges, with a relatively soft spectrum ( = 2.6). The half-year averaged photon flux then is one order of magnitude of that from the initial 9-year Fermi-LAT survey, indicating a strong -ray flare. B3 1239+376 is known as the unique blazar (candidate) in the direction and the association probability is calculated as 0.91, which establishes a physical connection between the -ray source and B3 1239+376. Moreover, by revisiting the Fermi-LAT data, prior -ray activities linking to the target are also found. Although these two detections are marginal (TS 15) each alone, a joint analysis brings enhanced statistics (TS = 31) enabling us to perform a localization analysis, from which B3 1239+376 also falls into the localization uncertainty, shown in Figure 2. Most importantly, behaviors of the -ray and infrared light curves are rather similar. At a scope of 16-year monitoring, -ray flux reaches up to a maximum level in year 2025, at the same time, the infrared fluxes are unprecedentedly high. In addition, the appearance of the prior tentative -ray signals is temporally coincident with the times of the high flux levels in the WISE light curves, see Figure 3. The coincidence acts as the decisive evidence to pin down the association relationship. The most energetic -ray photon relative to B3 1239+376 is at energy of 1Ā GeV, which could not challenge the current EBL models (e.g., Finke etĀ al., 2010). Although the data sampling of the WISE observations is limited, detections of contemporaneous brightening in the Fermi-LAT and infrared light curves play an important role in identification of high- -ray-emitting blazars (Liao etĀ al., 2019; Li etĀ al., 2020). It is also worth mentioning that the k-band core flux climbs when the -ray and infrared emissions brighten. Unfortunately, no continual VLBA monitoring after the activity is available and detecting a potential ejection of a new jet blob could not be accomplished then.
High-redshift (z 3) -ray blazars are rare. Among the thousands of extragalactic -ray emitters (Ajello etĀ al., 2022), only a handful of these sources are found (e.g., Ackermann etĀ al., 2017). Yet they are crucial for our understanding in AGN jet cosmic evolution, the star formation history of the universe, and the initial mass growth of SMBHs. In particular, besides the two sources above of 4, B3 1428+422 and GB 1508+5714 (Ackermann etĀ al., 2017; Liao etĀ al., 2018), there are five known blazars with 3.5 detected by Fermi-LAT to date. They are NVSS J121915+365718 ( = 3.53, Paliya etĀ al. 2020), PMN J2219-2719 ( = 3.63, Li etĀ al. 2020), PKS 0201+113 ( = 3.64, Lei etĀ al. 2024), MG3 J163554+3629 ( = 3.65, Ackermann etĀ al. 2017) and PKS 1351-018 ( = 3.72, Punsly etĀ al. 2021). Therefore, as an identified -ray emitter, B3 1239+376 ( = 3.82) is proposed as the third most distant -ray-emitting blazar so far. Most of these extreme sources are FSRQs with soft -ray spectra, but NVSS J121915+365718 is suggested as a BL Lac (Paliya etĀ al., 2020). B3 1239+376, along with other high- -ray sources, are plotted in a - diagram (i.e., Figure 6). The information listed in 4LAC-DR3 (Ajello etĀ al., 2022), as well as those reported in individual studies, are adopted. Considering the high apparent luminosity (up to in the half-year high flux state) and relatively soft spectrum (), it is not surprising that our target lies with the region occupied by the high- sources. In addition to the significant variations of the flux level, the spectral hardening when brighten is indicated for some sources. However, the investigation of spectral variation is not available for B3 1239+376 due to limited statistics in the first 12-year Fermi-LAT survey. Note that several distant -ray sources are not included in 4LAC-DR3, alike of B3 1239+376. Because the rather faint emissions, they are detectable only in the high flux state, while the background is overwhelming for the majority of the time. Therefore, a detailed temporal analysis is crucial to catch the -ray signal from high- blazars.
III Discussions and Summary
The multiwavelength observations not only bring valuable temporal information of B3 1239+376, but also allow us to draw the broadband SEDs, presented in Figure 7. According to the high flux state, quasi-simultaneous data including the -ray spectrum between MJD 58060 and 58242, ZTF g and r band photometry at MJD 58258 as well as the i band one at MJD 58257, WISE W1 and W2 band measurements at MJD 58260, and the RATAN-600 flux densities at MJD 58235 from 2.3 to 22Ā GHz. On the other hand, the low-flux-state SED consists of the Fermi-LAT upper limit from MJD 59065 to 60626, the Chandra snapshot at MJD 59944, the ZTF g band exposure at MJD 59937, and WISE W1 and W2 estimations at MJDĀ 59928. Additionally, -ray and infrared data correspond to the flare event in year 2025 are also presented, although they are not modeled due to the limited data coverage. Un-simulaneous archival data are collected and plotted as backgrounds, such as the radio data collected from NED, the optical photometric data from SDSS and multi-bands infrared data from Infrared Array Camera (IRAC) aboard Spitzer. No significant flux variations are seen in the optical r and g band, in which the accretion disk emission is likely dominant. The big blue bump is described as a standard Shakura & Sunyaev (1973) disk, extending from 3 to 2000, where is the Schwarzschild radius. The total disk luminosity is produced as , in which represents the accretion rate and the accretion efficiency is set as a typical value, 0.1. The accretion disk emission is characterized as a multi-temperature radial profile. At a certain radius , the local temperature is,
where is the Stefan-Boltzmann constant. Since the g band is likely drop-out due to neutral intergalactic medium absorption, it is not considered for the modeling of the disk. It is remarkable that the infrared radiations significantly vary for different flux states. In the low flux state, the W1 and W2 band data points are in alignment with the accretion disk radiations, while the fluxes increase, especially for the W2 band, in the high flux state when the jet contribution is un-ignorable. The jet activity in the infrared bands is also supported by the un-simulaneous IRAC observations (Teplitz etĀ al., 2010). The change of the -ray spectra is also clear. In the 4-year long low flux state, only an upper limit is yielded for Fermi-LAT. While a significant -ray signal are revealed in the half-year period of 2025. Due to the limited statistics, a butterfly plot is drawn for the -ray spectrum of the high-flux-state SED.
Here a classic homogeneous one-zone leptonic radiation scenario is used to give a description of the jet emission. A relativistically moving compact blob with a radius of embedded in the magnetic field (B) is responsible for the emission. Within the blob, it is assumed that the emitting electrons follow a broken power-law distribution,
where the are indices of the broken power-law distribution, , and are the minimum, breaking and maximum energies of the electrons, respectively, and is the normalization of the particle number density. The transformations of frequency and luminosity between different frames (i.e., the jet comoving frame and the observational frame) are and . In the leptonic radiation scenario, both synchrotron and IC processes are included. For the former, the synchrotron self-absorption (SSA) process is considered. In the IC processes, the KleinNishina effect is taken. Because of the strong Ly line (i.e., 2.5 Hz at the AGN frame) detected in the SDSS spectrum, it is adopted as the external soft photons in the EC process. Assuming the broad-line photons are 10% of the accretion luminosity in a shell-like shape located at a distance cm (e.g., Ghisellini etĀ al., 2011), the corresponding energy density is set as = 2 erg . The radius of the emitting blob is usually constrained by the observed variability timescale in short term, , however, no such information of B3 1239+376 is available. Thus, a typical value of cm is used.
The simple leptonic model can reasonably reproduce the both SEDs, see Figure 7. The corresponding input parameters are listed in Table 1. The radio data points are not considered in the jet modeling because they are likely from an extended region rather than the compact blob, and the radio emissions are seriously suffered by the SSA. Our jet modeling analyses suggest that B3 1239+376 is likely a low-synchrotron-peaked source (i.e., Hz, Padovani & Giommi 1995; Abdo etĀ al. 2010). In the high flux state, the synchrotron bump rises and shifts close to the infrared domain, resulting the increase of the flux densities there. On the other hand, the high energy bump moves into the detection energy range of Fermi-LAT at the same time, which leads to the emergence of the -ray signal. Theoretically, the injection of fresh emitting electrons in the dissipation region, as well as the enhancement of the magnetic field intensity, could account for the broadband variations. Therefore, it is not surprising that input parameters such as and B for the high flux state SED are larger than that of the other SED. The Compton dominance of B3 1239+376 is indicated as 10, consistent with that found in other high- blazars (e.g., Ghisellini etĀ al., 2011; Ackermann etĀ al., 2017). Interestingly, even in the low flux state, the X-ray emission of B3 1239+376 is luminous, which is also seen in several -ray quiet high- blazars (Marcotulli etĀ al., 2020, 2025). Together with the hard X-ray spectra, blazars are believed as powerful MeV beacons in the early cosmos. Meanwhile, an important jet parameter, the bulk Lorentz factor (), is worthy to be mentioned specially. In the classic orientation based AGN unified scheme, the number ratio of the jetted AGNs between the aligned and the apparent populations is suggested as . Thus, is a key parameter to determine the number density of jetted AGNs in the early cosmic time. The value for B3 1239+376 obtained here is 15, in agreement with the results (i. e., 1020) provided for other similar sources (e.g., Ghisellini etĀ al., 2011; Ackermann etĀ al., 2017; Liao etĀ al., 2020). In addition to the jet SED modeling studies, other approaches are used to probe the jet Lorentz factor of high- sources. Searching for the potential evidence of fast variability, especially in rays, can put a severe constraint on the value (e.g., Li etĀ al., 2018). Moreover, detections of the proper motion of the radio jets provide a direct estimation of the essential properties of the jet (e.g., Zhang etĀ al., 2020; Krezinger etĀ al., 2024). Considering the recent -ray activity of B3 1239+376 revealed by Fermi-LAT, quick follow-up radio interferometry observations that would reveal its jet kinematics are strongly encouraged.
In summary, based on detailed multi-wavelength investigations, B3 1239+376 ( = 3.82) is identified as a -ray emitter, and hence it is concluded as the third most distant -ray-emitting blazar so far, underneath B3 1428+422 ( = 4.72) and GB 1508+5714 ( = 4.3). Analysis of Fermi-LAT data yields a significantly detected (7.7 globally) -ray source cospatial with B3 1239+376 in a half-year period of 2025. It is known as the unique blazar (candidate) within the location uncertainty and an association probability of 0.91 is estimated, which suggest a physical connection between the -ray source and B3 1239+376. Meanwhile, indications of prior -ray activities are also found. Although for the two epochs each alone the statistics are limited, a joint analysis returns a boosted detection significance and the B3 1239+376 remains within the -ray localization error radius. Moreover, in a multi-wavelength perspective, the -ray and infrared fluxes contemporaneously brighten. In particular, emissions in these two windows reach up to the maximal level at the same time, under a total 16-year monitoring. The temporal coincidence helps us to determine the association relationship. Interestingly, accompanying radio core flux increase until Mar. 2018 is observed. In addition to the temporal behaviors, broadband SEDs of B3 1239+376 are plotted and described by a one-zone leptonic radiation scenario. Ejections of fresh emitting electrons into the dissipation region together with an increase of the magnetic field intensity may be responsible for the variations. Comparison between B3 1239+376 and other distant blazars are carried out. The soft -ray spectrum, luminous X-ray and -ray emissions, and variability across different bands, are in agreement with that found in other similar sources. To further investigate its jet properties, future multi-wavelength campaigns are urged.